# Offline Fictitious Self-Play for Competitive Games

### v1

@@ 4.3 Aggregate Average Strategy in Samples
Function Off-FictitiousSelfPlay($\mathcal{D}_{E}$):
      Estimate the empirical behavioural policy $\tilde{\pi}_{b}$$\Pi_{i}\leftarrow\{\pi_{b}^{i}\},\quad\pi_{0}\leftarrow\tilde{\pi}_{b}$$\mathcal{D}^{i}_{0}\leftarrow\{(s,a,\pi^{i}_{0}(s,a))|(s,a)\in\mathcal{F}^{i}(\mathcal{D}_{E})\}$for $k=1$ to $K$ do
           for each player $i\in\mathcal{N}$ do
                $\mathcal{D}^{i}_{k}\leftarrow$ GenerateDataset($\mathcal{D}_{E},\mathcal{D}^{-i}_{k-1}$) $\beta^{i}_{k}\leftarrow$ LearnBestResponse($\mathcal{D}^{i}_{k},\beta^{i}_{k-1}$) $\mathcal{D}^{i}_{k}\leftarrow$ UpdateAverageStrategy($\mathcal{D}^{i}_{k-1},\beta_{k}^{i}$) Update collection of BRs $\Pi_{i}\leftarrow\Pi_{i}\cup\{\beta_{k}^{i}\}$
     return Aggregate($\Pi$)
Function GenerateDataset($\mathcal{D}_{E},\mathcal{D}^{-i}_{k-1}$):
[…]
Function LearnBestResponse($\mathcal{D}^{i}_{k},\beta^{i}_{k-1}$):
      Initialize policy parameters with $\beta^{i}_{k}\leftarrow\beta^{i}_{k-1}$  Optimize $\beta^{i}_{k}$ with offline RL and data $\mathcal{D}^{i}_{k}$return $\beta^{i}_{k}$
Function UpdateAverageStrategy($\mathcal{D}^{i}_{k-1},\beta^{i}_{k}$):
      for $(s,a,\pi_{k-1}^{i}(s,a))\in\mathcal{D}^{i}_{k-1}$ do
           Calculate $\pi_{k}^{i}(s,a)$ with $\beta^{i}_{k},\pi^{i}_{k-1}$ following Equation 1, and update dataset $\mathcal{D}^{i}_{k}$
     return $\mathcal{D}^{i}_{k}$
4.4 Learning in Real-world Datasets
Considering an offline dataset sampled from extensive-form games, the two assumptions in Section 4.1 cannot be completely fulfilled. Initially, a discrepancy in the probability of trajectories is invariably present between game and sampled datasets. This results in an irremovable error in learning best responses. Leslie & Collins (2006) proposed weakened fictitious play that allows suboptimal $\epsilon$-best responses in the learning process. The discrepancy does not totally make fictitious play fail but affects the distance of the final solution from equilibrium. In practice, sampling a large dataset can mitigate the issue. The sample complexity is a common problem discussed in previous works about offline learning (Cui & Du, 2022a; Yan et al., 2022).
Secondly, it is not realistic to gather a fully covered dataset, especially in large-scale games. Previous works have discussed the requirements of the coverage to find NE (Cui & Du, 2022a; Li et al., 2022). However, their assumption is too restrictive to be fulfilled in real-world problems. In contrast, this work discusses how to minimize the distance to NE, i.e. the exploitability, under a given dataset, which is a more realistic setting. In fictitious self-play, a straightforward way to reduce exploitability is to learn better $\epsilon$-best responses at each step, which is the same for Off-FSP. The LearnBestResponse function can utilize any variants of the offline RL algorithms, which effectively address the issue of incomplete data set coverage.
Figure 4: Example partially covered datasets for Rock-Paper-Scissors. Numbers in the grids show the probability density of different samples. The red dashed boxes indicate the probability of different actions for corresponding behavioural strategies. (Left) A partially covered example with infinite samples. (Right) A partially covered example with finite samples.
Considering infinite samples, the missing of some $(s,a)$ is caused by $\pi_{b}(s,a)=0$ where $\pi_{b}$ is the behavior strategy to sample the dataset. We show this case in a simple matrix game, Rock-Paper-Scissors(RPS), with the first example of Figure 4. In this case, the covered part of the matrix can be seen as a subgame, where agents can only choose actions from observed ones. The payoff of the subgame is known. A conservative solution for this dataset is to find the Nash equilibrium of the subgame. However, if the lower bound of payoff in this game is known, allowing the selection of out-of-distribution actions may be a more reasonable setting. With the knowledge that payoffs of RPS are in the range of $[-1,1]$, player $1$ consistently selects paper as it yields a payoff of $1$. For player $2$, the only action, rock, in the dataset gets a payoff of $-1$, and choosing OOD actions may be a better selection because the rock is already the worst action. Some offline RL algorithms, such as CQL, give the OOD actions low rewards rather than set hard restrictions on these actions, which enables the learned policy to choose OOD actions. The experimental results in  Sections 5.1 and 5.2 show that restricting the policy space to subgames leads to higher exploitability.
[…]
5 Experiments
In this section, we design a series of experiments to show the performance of Off-FSP and investigate the OOD issue in offline competitive games. We first evaluate Off-FSP on toy examples in a matrix game, rock-paper-scissors. We compared Off-FSP against the state-of-the-art offline RL methods for competitive games, including OEF (Li et al., 2022), Behavior Cloning (BC), and other single-agent Offline RL methods, on an imperfect-information poker game, Leduc Hold’em. We also conduct ablation studies on Off-FSP in  Section D.2.
To demonstrate the compatibility of our approach with existing single-agent offline RL algorithms, we combine Off-FSP with three methods: CQL (Kumar et al., 2020), BCQ (Fujimoto et al., 2019b) and CRR (Wang et al., 2020). For variants with CQL, denoted as CQL-$\alpha$, we conduct different hyperparameters $\alpha$ which indicates the penalty level of choosing OOD actions. Higher $\alpha$ means less OOD actions performed by the policy. When $\alpha=0$, there is no penalty, and CQL degrades to DQN. More details about the experiments are provided in Appendix C.
5.1 Rock, Paper, Scissors
To provide empirical evidence for our statements in Section 4.4 and show the effectiveness of our methods, we first give results on three datasets of rock-paper-scissors. The first dataset, FULL, is randomly sampled from a uniform behavioural strategy, $\pi_{b}=(\frac{1}{3},\frac{1}{3},\frac{1}{3})$, which is a fully covered dataset and an approximation of ideal dataset. P1 and P2 are partially covered datasets described in Figure 4. Figure 5 shows the NashConv within $500$ iterations and the average strategies of player $2$ after the end of training.
In FULL dataset, both CQL-00 and CQL-$1$ with Off-FSP are converged to a $\epsilon$-NE with a small gap with online fictitious play. In the partially covered P2 dataset, only CQL with $\alpha>0$ can properly address the OOD issues successfully converges to a $\epsilon$-NE, while CQL-00 is misleading by one single sample of Rock2 and selects Rock2 with a probability of $1$. In the P1 scenario, employing CQL with a smaller $\alpha$ value results in a reduced NashConv. This outcome is attributed to the allowance of OOD actions, as elaborated in Section 4.4. The outcomes from experiments P1 and P2 demonstrate the complexities inherent in OOD problems. The findings suggest that neither arbitrary selection nor total exclusion of OOD actions is optimal, emphasizing the need for sophisticated offline algorithms.
Figure 5: Results of RPS. (Left) NashConv. (Right) Player $2$’s Average Strategy.
5.2 Leduc Hold’em
(a) NashConv
(b) OOD ratio
(c) NashConv w/o OOD actions
(d) Payoff
Figure 6: Results of Leduc Hold’em
We now show the results compared with existing algorithms for offline competitive games. We collect 5 datasets on Leduc Hold’em with different quality. Mix and Random are nearly fully covered datasets, which are sampled by a random policy and aggregate strategy with expert and random policy. P20, P10, and P5 are partially covered datasets sampled from mixed strategies of $20,10,5$ best responses. These best responses are randomly picked from the training process of online PSRO. The more BR picked for sampling, the more state-action pairs are covered by the datasets. We visualize the ratio of covered state-action pair for all datasets in Appendix D. All these datasets contain 10 0001000010\,00010 000 trajectories. Mix and P20 datasets sampled from a high-quality strategy, so we can easily obtain a good strategy by supervised learning their empirical behavioral strategies $\tilde{\pi}_{b}$. This paradigm is also called behavior cloning, which is a popular baseline for offline learning.
We evaluate state-of-the-art offline learning baselines and variants of Off-FSP on all the datasets. Figure 5(a) shows the NashConv of all the algorithms, and all variants of our method outperform other baselines. The result also demonstrates the ability of Off-FSP to robustly integrate with any existing offline RL methods. BC supervised learning the empirical behavioral strategies $\tilde{\pi}_{b}$ of all datasets, which shows Off-FSP can also outperform the sampling strategy. Mix and P20 datasets are both sampled from high-quality strategies, while the Random dataset is sampled from a random strategy. OEF-DCFR and OEF-PSRO are two variants of OEF (Li et al., 2022), which is a model-based offline algorithm for finding NE. OEF mixed its strategy with BC, evaluated it in online games multiple times and used the minimum NashConv as its results. This operation involves interaction with online games, which violates the offline scenario. To fair comparison, we remove the mixing operation in our main experiment. The results with mixing are shown in  Appendix D. OEF operates based on the assumption of full data coverage, so OEF strategies have a high probability of performing OOD actions on partially covered datasets, as shown in Figure 5(b).
To investigate the influence of OOD actions, we restrict all the strategies on the observed state-action pairs and evaluate these restricted strategies. Figure 5(c) demonstrates significant performance degradation in some methods. We also learned an online CFR to find a $\epsilon$-NE on the subgames of different datasets. For partially covered datasets, the $\epsilon$-NE shows large NashConv values. This indicates that directly solving NE in subgames is insufficient for resolving the offline learning problem in competitive games.
In previous works (Mathieu et al., 2023; Qu et al., 2023), directly applying single-agent Offline RL on the fixed dataset is also an attempt to solve offline competitive games. This paradigm is identical to learning best responses to the behavioural strategies $\pi_{b}$. These strategies overfit their opponents and are vulnerable to exploitation. These works only pay attention to strategies with higher returns and do not discuss NE. Therefore, we show the payoff of variants of Off-FSP playing with corresponding single-agent Offline RL algorithms, as Figure 5(d). Our methods show positive payoff to all the single-agent methods.
6 Conclusion
In this paper, we study offline multi-agent reinforcement learning for competitive games and propose Off-SP and Off-FSP to enable single-agent offline RL algorithms to be applied in this scenario. We find that Off-FSP can approximate NE even with partially covered datasets. Extensive experiments show that all variants of Off-FSP significantly outperform state-of-the-art baselines in different datasets of an imperfect-information two-player zero-sum game, Leduc Hold’em.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Derivation of Theorem
@@
Now we derive the weight when changing the opponent to $\pi_{-i}$. Since we do not need to adjust player $i$’s strategy, player $i$’s strategy is still considered as $\tilde{\pi}_{b}^{i}$. In player $i$’s perspective, the reach probability of a tuple $d_{t}=(s^{i}_{t},a^{i}_{t},s^{i}_{t+1})$ is
$Pr(d_{t}|\pi^{-i})=\Pi_{j=0}^{t}\pi^{i}(s^{i}_{j},a^{i}_{j})\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})\cdot\pi^{i}(a^{i}_{t+1}|s^{i}_{t+1}),$
@@
$=\frac{\Pi_{j=0}^{t}\tilde{\pi}_{b}^{i}(s^{i}_{j},a^{i}_{j})\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})}{\Pi_{j=0}^{t}\tilde{\pi}_{b}^{i}(s^{i}_{j},a^{i}_{j})\mathcal{T}_{\tilde{\pi}_{b}^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})}\cdot\frac{\tilde{\pi}_{b}^{i}(a^{i}_{t+1}|s^{i}_{t+1})}{\tilde{\pi}_{b}^{i}(a^{i}_{t+1}|s^{i}_{t+1})}$
@@
Considering $\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})$ and $\mathcal{T}_{\tilde{\pi}_{b}^{-i}}(s^{i}_{j+1}|s_{j},a_{j})$, both of these two items are the multiplication of a series of transition probabilities, of which only the probability of opponents’ strategies $\pi^{-i}$ is different. So the above equation can be simplified to:
@@ Appendix B Baselines
Conservative Q-Learning (Kumar et al., 2020) learns a conservative Q-function such that the expected value of a policy under this Q-function lower-bounds its actual value in order to alleviate overestimation of values induced by the distributional shift between dataset and the learned policy. In practice, we solve such an optimization problem to update our Q-function, where the first term is the penalty for lower-bounding the Q-function, and the second term is the Bellman error. We use CQL based on the QRDQN.
@@
Critic Regularized Regression (Wang et al., 2020) handles the problem of offline policy optimization by value-filtered regression. It selectively imitates the dataset by choosing a function $f$ that is monotonically increasing in $Q_{\theta}$.
[…]
@@
Batch-Constrained deep Q-learning (Fujimoto et al., 2019b) uses a generative model to learn the distribution of transitions in the dataset, samples several actions from it in the current state, perturbates them and chooses the one with highest Q value. In this way, they tried to minimize the distance of selected actions to the dataset and lead to states similar to those in the dataset. In the experiments, we use a discrete BCQ variant introduced in (Fujimoto et al., 2019a).
@@
Offline Equilibrium Finding (Li et al., 2022) is a model-based framework to find the game equilibrium in the offline setting. They directly train a model to describe the game dynamics and apply online equilibrium-finding algorithms like PSRO and DCFR (a deep learning version of CFR) to compute equilibrium. They also combine the behavior cloning policy with the model-based policy for improvement, where they mix these two policies with different weights and evaluate each weight in the online game to find the best combination.
Appendix C Details of experiments
C.1 Datasets
C.1.1 Rock, Paper, Scissors
FULL is randomly sampled from a uniform behavioral strategy, $\pi_{b}=(\frac{1}{3},\frac{1}{3},\frac{1}{3})$, i.e., every player randomly picks each action with the same probability. P1 and P2 are generated according to the exact proportion described in Figure 4, e.g., P1 comprises 500 episodes of Rock to Rock and 500 episodes of Paper to Rock.
Every dataset is composed of 1000 trajectories.
C.1.2 Leduc Hold’em
Random is sampled by uniformly randomly picking action from all legal actions in every decision node.
Mix is sampled from a mixed strategy, which is a 3-to-1 mix of expert and random strategies. The expert strategy is trained by PSRO with 30 iterations and is a mix of 30 deterministic policies according to a specific ratio.
Partial dataset (5-10-20) is sampled by first picking 5, 10, and 20 policies from the 30 components of the expert strategy and then uniformly mixing them to form the sampling strategy.
Every dataset comprises 10000 trajectories except those used in ablation study about the dataset size.
C.2 Variants of Off-FSP
@@
Off-FSP-CQl, Off-FSP-CRR, and Off-FSP-CQL are different variants of our methods that integrate with different single-agent offline RL methods in the function of LearnBestResponse. We then introduce details about these Offline RL methods.
@@
In CQL, we use a QRDQN as the base algorithm and a simple MLP as the Q network. Hence, the network’s input is the observation vector, and the output has a dimension of the number of quantiles times the number of actions. The hidden layer of MLP contains the same number of neurons. We use ReLU as the activation function between every linear layer.
@@
In CRR, the actor and critic networks have the same structure, first a shared feature network and then a MLP head. The shared feature network is a MLP with two layers, and the size of the second layer is the number of actions. All activation functions are ReLU. CQL penalty term is also added to the CRR objective function for better performance.
@@
In BCQ, the policy network and the imitation network have the same structure, which is also first a shared feature network and then an MLP head, exactly the same as CRR.
C.3 Hyperparameters
In OEF, all settings are the same as Li et al. (2022).
In RPS, we use Adam optimizer with a learning rate of 0.005 and betas (0.9, 0.999), and the network has only one layer, and the number of quantiles in QRDQN is 100. The target network is not used. In each iteration, the network is trained five epochs and is updated 100 times each epoch with a batch size of 1000.
In Leduc, the hyperparamters are shown in Table 1. All networks are optimized by Adam Optimizer (Kingma & Ba, 2017). The hidden layer number of CRR and BCQ describes the whole network, including the feature network and MLP head. The target network is used.
Table 1: Hyperparameters used in Leduc
@@ C.3 Hyperparameters
Training steps per iteration $M$ 1 000 1 000 1 000
@@
Appendix D Supplement Results
D.1 Analysis of Datasets
Figure 7: State-Action Cover Ratio
We visualize the state-action cover ratio of different datasets in Figure 7
D.2 Ablations
(a) Value of $\alpha$
(b) Size of Data
Figure 8: Results of ablations.
[…]
D.3 Learning Curves of Off-FSP
OEF evaluates strategies’ performance by allowing strategies to mix with BC by different weights. However, the results are decided by the best-mixed strategies. This is equivalent to allowing strategies to be interactively evaluated online, which is contrary to the setting of offline scenarios. And we cannot tell whether this result is caused by the original strategy or BC. We still provide the final results under this evaluation method in Figure 9, and our method still outperforms OEF.
Figure 9: NashConv with evaluation method in OEF

### v2

@@ 36  Optimize $\beta^{i}_{k}$ with offline RL and data $\mathcal{D}^{i}_{k}$;
Appendix B Related Work
In competitive games, also known as zero-sum games, people have developed a series of learning algorithms, where finding Nash equilibrium with self-play is the major learning paradigm for this problem. Fictitious self-play (FSP) (Heinrich, Lanctot, and Silver 2015) combines fictitious play (FP) (Brown 1951) with self-play and provably converges to an NE in extensive-form games. Counterfactual regret minimization (CFR) (Bowling et al. 2015) combines regret minimization with self-play, first solving an imperfect-information game of real-world scale. DeepNash (Perolat et al. 2022) extends regularised Nash dynamics with RL and converges to an approximate NE in Stratego. Policy-Space Response Oracles (PSRO) (Lanctot et al. 2017) learn to find NE by iteratively learning best responses to its policies’ population, which can also be seen as a population-based self-play. Some works (Silver et al. 2017; Ye et al. 2020; Yang et al. 2022) of large-scale games aim for a higher winning percentage or returns and try to exploit opponents, not pursuing strict NE, still based on the self-play paradigm.
In the offline learning paradigm, behavior cloning (BC) is the simplest form that directly imitates the sampling policy of datasets. The performance of BC is limited by the sampling policy in the dataset, while offline RL can surpass this limitation by maximizing returns. Offline RL faces the challenge of the extrapolation error on out-of-distribution states and actions. Batch-constrained deep q-learning (BCQ) (Fujimoto, Meger, and Precup 2019) and conservative q-learning (CQL) (Kumar et al. 2020) mitigate this issue by constraining the gap between policy and data distribution. Critic regularized regression (CRR) (Wang et al. 2020) solve this problem by weighted behavior cloning.
In offline competitive scenarios, BC is a viable algorithm, still limited by the quality of sampling policy in datasets. Mathieu et al. (2023); Qu et al. (2023) directly learn policies by single-agent offline RL without self-play. This way also requires high-quality behavior policies in datasets and is easily overfitted to fixed opponents. Cui and Du (2022b); Zhong et al. (2022); Cui and Du (2022a) propose theoretically feasible algorithms. Li et al. (2022) proposed a model-based learning framework, OEF, for extending online equilibrium-finding algorithms in offline scenarios. However, OEF only supports datasets with full coverage of state-action spaces, which is not realistic for real-world problems. Although previous theoretic methods allow for partially covered datasets, they cannot be practically applied to real-life problems. In this paper, Off-SP enables the self-play paradigm in offline MARL of competitive games. Off-FSP implement a practical algorithm to find NE and has the potential to learn on partially covered datasets by integrating any variants of single-agent offline RL algorithms.
Appendix C Derivation of Theorem
@@
Now we derive the weight when changing the opponent to $\pi_{-i}$. Since we do not need to adjust player $i$’s policy, player $i$’s policy is still considered as $\tilde{\pi}_{b}^{i}$. In player $i$’s perspective, the reach probability of a tuple $d_{t}=(s^{i}_{t},a^{i}_{t},s^{i}_{t+1})$ is
$Pr(d_{t}|\pi^{-i})=Pr(s^{i}_{0})\pi^{i}(a^{i}_{t+1}|s^{i}_{t+1})\cdot\Pi_{j=0}^{t}\pi^{i}(s^{i}_{j},a^{i}_{j})\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j}),$
@@
$=\frac{Pr(s^{i}_{0})\tilde{\pi}_{b}^{i}(a^{i}_{t+1}|s^{i}_{t+1})}{Pr(s^{i}_{0})\tilde{\pi}_{b}^{i}(a^{i}_{t+1}|s^{i}_{t+1})}\cdot\frac{\Pi_{j=0}^{t}\tilde{\pi}_{b}^{i}(s^{i}_{j},a^{i}_{j})\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})}{\Pi_{j=0}^{t}\tilde{\pi}_{b}^{i}(s^{i}_{j},a^{i}_{j})\mathcal{T}_{\tilde{\pi}_{b}^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})}$
@@
Considering $\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})$ and $\mathcal{T}_{\tilde{\pi}_{b}^{-i}}(s^{i}_{j+1}|s_{j},a_{j})$, both of these two items are the multiplication of a series of transition probabilities, of which only the probability of opponents’ policies $\pi^{-i}$ is different. So the above equation can be simplified to:
@@ Appendix D Baselines
Conservative Q-Learning (Kumar et al. 2020) learns a conservative Q-function such that the expected value of a policy under this Q-function lower-bounds its actual value in order to alleviate overestimation of values induced by the distributional shift between the dataset and the learned policy. In practice, we solve such an optimization problem to update our Q-function, where the first term is the penalty for lower-bounding the Q-function, and the second term is the Bellman error. We use CQL based on the QRDQN.
@@
Critic Regularized Regression (Wang et al. 2020) handles the problem of offline policy optimization by value-filtered regression. It selectively imitates the dataset by choosing a function $f$ that is monotonically increasing in $Q_{\theta}$.
$\underset{\pi}{\arg\max}\ \mathbb{E}_{(\mathbf{s},\mathbf{a})\sim\mathcal{D}}\left[f\left(Q_{\theta},\pi,\mathbf{s},\mathbf{a}\right)\log\pi(\mathbf{a}\mid\mathbf{s})\right]$ (5)
@@
Batch-Constrained deep Q-learning (Fujimoto, Meger, and Precup 2019) uses a generative model to learn the distribution of transitions in the dataset, samples several actions from it in the current state, perturbates them, and chooses the one with the highest Q value. In this way, they tried to minimize the distance of selected actions to the dataset and lead to states similar to those in the dataset. In the experiments, we use a discrete BCQ variant introduced in (Fujimoto et al. 2019).
@@
Offline Equilibrium Finding (Li et al. 2022) is a model-based framework to find the game equilibrium in the offline setting. They directly train a model to describe the game dynamics and apply online equilibrium-finding algorithms like PSRO and DCFR (a deep learning version of CFR) to compute equilibrium. They also combine the behavior cloning policy with the model-based policy for improvement, where they mix these two policies with different weights and evaluate each weight in the online game to find the best combination.
Appendix E Details of experiments
Datasets
Datasets of Rock, Paper, Scissors
D1 is randomly sampled from a policy, $\pi_{b}=(0.6,0.2,0.2)$. D2 is generated according to the exact proportion described in Figure˜3. Every dataset is composed of 1000 trajectories.
Datasets of Extensive-form Games
The Nash Conv of expert policies are $0.86,0.26,1.01$ in Leduc Poker, Large Kuhn Poker, and Oshi Zumo, separately. Every dataset comprises 10 000 trajectories The size of samples in datasets is far less than the number of samples required by online algorithms. Taking PSRO as an example, training BR requires about 10 000 trajectories at each iteration, and additional samples are needed to evaluate the new BR.
Rock, Paper, Scissors
As described in Figure˜3, the first dataset, D1, is randomly sampled from a uniform policy, $\pi_{b}=(0.6,0.2,0.2)$, which is a fully covered dataset and an approximation of the ideal dataset. D2 are partially covered datasets, which are constructed following the probability. Figure˜3 shows the NashConv within $500$ iterations and the average policies of player $2$ after the end of training.
In D1 dataset, BC learns a suboptimal policy with an exploitability of $0.4$, while the exploitability of Off-FSP is less than $0.1$ after $500$ steps. Off-FSP-DQN is the ablation variant of Off-FSP without the offline RL algorithm, to learn the best response. It is easy to be misled by the OOD actions in D2. Off-FSP-CQL is the default setting of Off-FSP, which successfully learns a robust policy with low exploitability in both D1 and D2 datasets.
Extensive-form Games
Leduc Poker is a simplified version of two-player Texas holdem. Leduc Poker is a widely-used imperfect information normal-form game as a testbed for online algorithms.
Large Kuhn Poker is a variant of Kuhn Poker where an initial pot for each player is $5$. Each player has four actions: Fold, Check, Call, and Raise $1$ pot. Only in the first $8$ steps can players raise.
Oshi-Zumo is a deterministic board game in which two players repeatedly bid to push a token off the other side of the board (Buro 2004). The size of the board is $3$. The initial number of coins for each player is $4$. The maximum game horizon is $6$.
Ball Defence Game
The robot arm in the ball defence game is Agilex PiPER, a 6-DOF arm. It is equipped with an RGBD camera, Zed mini, to perceive the position of the blue ball. We use color segmentation with 1344x376 resolution and 100 FPS to detect the ball. The size of each basket, i.e., the goals, is 20x17x10 cm, and the size of the black racket is 16x13 cm.
In data collection, we restricted the action space of the robotic arm’s end effector to the area above the basket and discretized it into a 10-dimensional discrete space. This increases the success rate of data collection and ensures that the robot’s movements will not cause any damage to the experimental setup.
In the evaluation phase, we invited 10 human players to play the ball defence game and evaluate the performance of all methods. For fair comparison, each human player first has a 3-minute warmup to familiarize themselves with this game, and then plays with each method in a random order. For each method, we only retain the first two results that each person throws at each goal. The total number of evaluations for each goal and each method is 20.
Variants of Off-FSP
@@ Appendix E Details of experiments
Off-FSP-CQL, Off-FSP-CRR, and Off-FSP-BCQ are different variants of our methods that integrate with different single-agent offline RL methods in the function of LearnBestResponse. We then introduce details about these Offline RL methods.
@@
In CQL, we use a QRDQN as the base algorithm and a simple MLP as the Q network. Hence, the network’s input is the observation vector, and the output has a dimension of the number of quantiles, multiplied by the number of actions. The hidden layer of MLP contains the same number of neurons. We use ReLU as the activation function between every linear layer.
@@
In CRR, the actor and critic networks have the same structure, first a shared feature network, and then an MLP head. The shared feature network is an MLP with two layers, and the size of the second layer is the number of actions. All activation functions are ReLU. The CQL penalty term is also added to the CRR objective function for better performance.
@@
In BCQ, the policy network and the imitation network have the same structure. They are also first a shared feature network and then an MLP head, which is exactly the same as CRR.
Table 2: Hyperparameters used in Leduc Poker
@@
Training iterations 2 000 2 000 2 000
Training steps per iteration 1 000 1 000 1 000
@@
Table 3: Hyperparameters used in Large Kuhn Poker, Oshi Zumo and ball Defence Game.
Hyperparameter Large Kuhn Poker Oshi Zumo Ball Defence
Offline RL algorithm CQL CQL CQL
Hidden layer number 4 4 4
Hidden layer size 256 256 256
Learning rate 0.0001 0.0001 0.0001
Adam $(\beta_{1},\beta_{2})$ (0.9, 0.999) (0.9, 0.999) (0.9, 0.999)
Training iterations 1 000 1 000 200
Training steps per iteration 1 000 1 000 50
Target Network update frequency 100 100 10
Batch size 1024 1024 128
QRDQN number of quantiles 100 100 100
CQL $\alpha$ 0.1 0.01 0.1
Hyperparameters
For OEF baseline, all settings are the same as Li et al. (2022).
In RPS, we use Adam optimizer with a learning rate of 0.005 and betas (0.9, 0.999), and the network has only one layer, and the number of quantiles in QRDQN is 100. The target network is not used. In each iteration, the network is trained for five epochs and is updated 100 times each epoch with a batch size of 1000.
In extensive-form games, the hyperparameters are shown in Tables˜2 and 3. All networks are optimized by Adam Optimizer (Kingma and Ba 2017). The number of hidden layers in CRR and BCQ describes the whole network, including the feature network and MLP head. The target network is used. Our experiments run on AMD EPYC 7302 CPU and 3080Ti GPUs.
In the ball defence game, the hyperparameters are shown in Table˜3. To accelerate the training process, we use softmax with temperature 0.5 to calculate the probabilities of each action. In the final evaluation, we only keep a checkpoint of the robot policy at 100, 125, 150, 175, and 200 iterations and randomly pick one of them each time.
Appendix F Explanation on RPS experiments
As described in Figure˜3, the first dataset, D1, is randomly sampled from a uniform behavioural strategy, $\pi_{b}=(0.6,0.2,0.2)$, which is a fully covered dataset and an approximation of the ideal dataset. D2 are partially covered datasets, which are constructed following the probability. Figure˜3 shows the NashConv within $500$ iterations and the average strategies of player $2$ after the end of training.
In D1 dataset, BC learns a suboptimal strategy with an exploitability of $0.4$, while the exploitability of Off-FSP is less than $0.1$ after $500$ steps. Off-FSP-DQN is the variant of Off-FSP by using DQN, an online RL algorithm, to learn the best response. The algorithm only works well in ideal datasets such as D1, and it is easy to be misled by the OOD actions in D2. Off-FSP-CQL is the default setting of Off-FSP, which successfully learns a robust strategy with low exploitability in both D1 and D2 datasets.
Appendix G Supplement Results
Learning Curves of Nash Conv for Extensive-form Games
Figure 10: Learning Curves of our main experiments.
Figure˜10 shows all the Nash Conv’s learning curves of the main experiments in Section˜5.1. CQL is Q-learning algorithms with Q function only.
Figure 11: Results on Extensive-Form Games with evaluation method in OEF. (Top) NashConv on Mix Datasets; (Bottom) NashConv on Population Datasets.
Experimental Results Under Rules of OEF
OEF evaluates policies’ performance by allowing policies to mix with BC by different weights. However, the results are decided by the best-mixed policies. This is equivalent to allowing policies to be interactively evaluated online, which is contrary to the setting of offline scenarios. It is difficult to figure out whether the original policy or BC causes this result. We still provide the final results under this evaluation method in Figure˜11, and our method still outperforms OEF.